# The morphology and interface structure of titanium on graphene

### v1

@@ VII.6 Atomic force microscopy
VII.7 Winterbottom reconstruction calculations
@@
VII.8 Density Functional Theory Calculations
@@
The DFT calculations are carried out using VASP 5.2 [74] employing the Perdew, Burke, and Ernzerhof (PBE) Generalized Gradient Approximation (GGA) [75] for the exchange-correlation function and PAW pseudopotentials [76, 77]. Van der Waals interactions are accounted for using the Tkatchenko-Scheffler correction [78, 79]. The electronic wavefunctions are expanded with planewaves with energy less than 400 eV. The electronic self-consistent loop is terminated when energy changes are less than $1\cdot 10^{-6}$ eV and the ionic relaxations are considered converged when the magnitude of the largest force on any atom is less than 0.01 eV/Å. The Gr slab is modeled using a 3$\times$3 surface supercell created from graphite with hexagonal lattice with lattice constants 2.458 and 6.592 Å. To mitigate the fictitious interactions between the images in the non-periodic direction perpendicular to the Gr slab, we used a supercell approach with at least 12 Å of vacuum. The Brillouin zone is sampled using the $\times 5\times$1 Monkhorst-Pack grid. All of the calculations are spin-polarized. To determine the interface, we used the Zur and McGill [80] lattice matching scheme in conjunction with lattice shifs [81, 82]. The lattice matching scheme accounts for possible rotations of one lattice with respect to the other, or expansion of both lattices, in addition to lattice reduction scheme, to minimize the strain and lattice mismatch between the two lattices. The precision of the match gives an upper bound on the strain required to form chemical bonds across the interface. The lattice shifts ensure that the atom registry at the interface is optimum [81].
VIII Acknowledgements
@@
This work is primarily supported through the U.S. Department of Energy, Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division under FWP ERKCK47 ‘Understanding and Controlling Entangled and Correlated Quantum States in Confined Solid-state Systems Created via Atomic Scale Manipulation’, as well as partially supported by the Quantum Science Center (QSC), a National Quantum Information Research Center of the U.S. Department of Energy (DOE) on probing quantum matter. W.A.S. is partially supported by the US National Science Foundation (Award No. CBET-2130804). Computational support was provided in part by the University of Pittsburgh Center for Research Computing through the resources provided on the H2P cluster, which is supported by NSF (Award No. OAC-2117681). J.D.T. acknowledges support from Independent Research Fund Denmark through grant no. 9035-00006B. K.R. acknowledges funding from the Hugh Hampton Young Memorial Fund at MIT. J.J.P. acknowledges funding from the Natural Sciences and Engineering Research Council of Canada and MIT School of Engineering. The authors would like to acknowledge Michael Tarkanian for help in manufacturing TEM sample holders and Prof. Jeehwan Kim for equipment access. This work made use of the Materials Research Laboratory Shared Experimental Facilities at MIT, supported in part by the MRSEC Program of the National Science Foundation under Award No. DMR1419807. This work was carried out in part using MIT.nano’s facilities. This work was performed in part at the Center for Nanoscale Systems (CNS), a member of the National Nanotechnology Coordinated Infrastructure Network (NNCI), which was supported by the National Science Foundation under NSF Award No. 1541959. Figure 1(d) and 2(c) were created using Vesta [83].

### v2

@@ III.6 Atomic force microscopy
III.7 Winterbottom reconstruction calculations
@@
III.8 Density Functional Theory Calculations
@@
The DFT calculations are carried out using VASP 5.2 [80] employing the Perdew, Burke, and Ernzerhof (PBE) Generalized Gradient Approximation (GGA) [81] for the exchange-correlation function and PAW pseudopotentials [82, 83]. Van der Waals interactions are accounted for using the Tkatchenko-Scheffler correction [84, 85]. The electronic wavefunctions are expanded with planewaves with energy less than 400 eV. The electronic self-consistent loop is terminated when energy changes are less than $1\cdot 10^{-6}$ eV and the ionic relaxations are considered converged when the magnitude of the largest force on any atom is less than 0.01 eV/Å. The Gr slab is modeled using a 3$\times$3 surface supercell created from graphite with hexagonal lattice constants 2.458 and 6.592 Å. For the Ti layers, we used a slab comprising of 15 layers, where the top 6 layers (those furthest from the Gr) were fixed at their bulk positions to minimize finite-size effects. For simplicity, Fig. 1(f) shows only the first 4 Ti layers. To mitigate the fictitious interactions between the images in the non-periodic direction perpendicular to the Gr slab, we used a supercell approach with at least 12 Å of vacuum. The Brillouin zone is sampled using the $\times 5\times$1 Monkhorst-Pack grid. All of the calculations are spin-polarized. To determine the interface structure, we used the Zur and McGill [86] lattice matching scheme in conjunction with lattice shifs [87, 88]. The lattice matching scheme accounts for possible rotations of one lattice with respect to the other, or expansion of both lattices, in addition to the lattice reduction scheme, to minimize the strain and lattice mismatch between the two lattices. The precision of the match gives an upper bound on the strain required to form chemical bonds across the interface. The lattice shifts ensure that the atom registry at the interface is optimum [87].
IV Supporting Information
The Supporting Information is available free of charge at [insert link here].
Optical microscopy sample overview images (Fig. S1), high-magnification TEM images of 1L Gr with Ti deposited at room temperature (Fig. S2), nucleation density measurement data at room temperature (Fig. S3), additional TEM data for Ti depositions at elevated temperature including nucleation density measurements and high-magnification images (Fig. S4-S8), DF-TEM images of 1L Gr with Ti deposited at 400 °$\degree$°C, Raman spectroscopy data obtained during sample fabrication (Fig. 10), additional DFT results (Table S1), scattering data from x-ray scattering for Gr, Ti, TiC, and TiO2 (Table S2).
V Acknowledgements
@@
This work is primarily supported through the U.S. Department of Energy, Office of Science, Basic Energy Sciences (BES), Materials Sciences and Engineering Division under FWP ERKCK47 ‘Understanding and Controlling Entangled and Correlated Quantum States in Confined Solid-state Systems Created via Atomic Scale Manipulation’, as well as partially supported by the Quantum Science Center (QSC), a National Quantum Information Research Center of the U.S. Department of Energy (DOE) on probing quantum matter. W.A.S. is partially supported by the US National Science Foundation (award no. CBET-2130804). Computational support was provided by the University of Pittsburgh Center for Research Computing through the resources provided on the H2P cluster, which is supported by NSF (award no. OAC-2117681). J.D.T. acknowledges support from Independent Research Fund Denmark through grant no. 9035-00006B. K.R. acknowledges funding from the Hugh Hampton Young Memorial Fund at MIT. J.J.P. acknowledges funding from the Natural Sciences and Engineering Research Council of Canada and MIT School of Engineering. The authors would like to acknowledge Michael Tarkanian for help in manufacturing TEM sample holders and Prof. Jeehwan Kim for equipment access. This work made use of the Materials Research Laboratory Shared Experimental Facilities at MIT, supported in part by the MRSEC Program of the National Science Foundation under award no. DMR1419807. This work was carried out in part using MIT.nano’s facilities. Figure 1(d) and 2(c) were created using Vesta [89].